Ahh...please help. How do I set this up??? That's all I need. I am confused.
A man goes into a grocery store and buys a grapefruit and a toy ball. A 7% tax is applied to the price of non food items. If the total bill is $5.15, including $.018 in tax on the toy ball, what are the prices of the grapefruit and the ball? Thanks.. :)

Answer by ikdeep(189)

Man purchased two items i.e grapefruit and a toy ball..
He gave 7% tax on non food item and we know that a toy ball is a non food item .
So, he gave 7% tax on toy ball..
Now, we are also given that tax on toy ball = $0.018
Therefore 7% tax on toy ball = $ 0.018,,,
let price toy ball be x........so we get
7% of x = $ 0.018
or 7/100 * x = 0.018
0.07 * x = 0.018
x = 0.018 / 0.07
x = $0.26 approximately
or price of toy ball = $0.26
Now , we are also given that his total bill = $5.15
Therefore, Price of grape fruit + price of toy ball = $ 5.15
Price of grape fruit + $0.26 = $ 5.15
Price of grape fruit = $ 5.15 - $0.26
Price of grape fruit = $ 5.15 - $0.26
= $4.89


Please feel free to revert back for any further queries.